OpenStudy (anonymous):

2/3t + v = r solve for t

OpenStudy (anonymous):

2/3t + v = r subtract v from both sides: 2/3t + v - v= r - v 2/3t = r - v multiply 3/2 to both sides: (3/2)(2/3t) = 3/2(r - v) t = 3/2(r - v)